# Unitary quantum gravitational physics and the CMB parity asymmetry

### v1

@@ Abstract
The proposal of Direct-Sum Quantum Field Theory (DQFT) offers a new perspective for quantum fields by combining parity and time reversal operations, blurring the distinction between quantum past and future while preserving causality. This approach provides a unitary QFT in curved spacetime, resolving the information-loss paradox. When applied to inflationary quantum fluctuations, DQFT predicts variations in CMB measurements, explaining longstanding anomalies as a result of parity asymmetry. The data strongly supports the unitary treatment of quantum fluctuations, with a probability over 650 times greater than the standard prediction. This significant discrepancy underscores the validity and importance of the DQFT approach in understanding the intricate relationship between gravity and quantum mechanics.
@@
Introduction - Understanding and testing the interplay between gravity and quantum mechanics (QM) is the most important key to unraveling the new physics beyond the standard model of particle physics and Einstein’s general relativity (GR). The merge of GR and QM is an essential physics to be unlocked at the length scales of gravitational horizons. For this, we perhaps need an in-depth understanding of how one can resolve an age-old conundrum in the concept of time between the GR and QM. The success of Quantum field theory (QFT) in Minkowski spacetime Coleman (2018) already teaches us several lessons in this regard. Indeed QFT is the culmination of special relativity (SR) and QM where the concept of time already enters into the conflict. In SR, the space and time are on equal footing but in QM spatial coordinates are elevated to operators while time is just a parameter due to its anti-unitary nature under discrete transformation ($\mathcal{T}:t\to-t$). The SR and QM are merged by imposing the commutativity of operators corresponding to space-like distances. The underlying reason is that SR forbids any sort of communication happening in space-like distances and we respect it and achieve it in formulating the rules of second quantization. This is exactly the lesson one must keep in mind for understanding quantum fields in curved spacetime. That is we take lessons from classical physics and apply them according to the rules of quantum physics. The quantum field theory in curved spacetime (QFTCS) is the first step in seeing GR and QM together besides the renormalizable ultra-violet complete quantum gravity that may be required at the Planck scales. Furthermore, QFTCS is a necessity in the context of understanding inflationary quantum fluctuations Mukhanov and Chibisov (1981); Sasaki (1986).
@@
In this letter, we address crucial questions related to the meaning of time reversal and parity inversions in the context of inflationary background spacetime to place quantum fields on it and uncover their true nature in the cosmic microwave background (CMB). We explain how the direct-sum quantum field theory (DQFT) is a necessity to resolve the unitarity and information-loss problem of QFTCS. The question of unitarity in QFTCS precedes the question of a completely renormalizable quantum gravity theory of Planck scales. In contrast to the widespread schools of thought that argue unitarity in QFTCS needs to be given up within GR and QM, we came up with a fundamental understanding of it and an observational test of it by the remarkable revelation of parity asymmetry in the CMB. We find that the theory of inflationary quantum fluctuations with DQFT (which we call shortly DSI which stands for ”Direct-Sum inflation”) fits up to 650 times better than the Standard Inflation (SI). This letter ultimately presents the picture of how a new theory of QFTCS emerged from fundamental questions of quantum gravitational physics that can potentially explain the observed CMB large-scale anomalies. The new results in this letter are further complemented with our companion paper Gaztañaga and Sravan Kumar (2024).
@@
DQFT, in a nutshell, ($\hbar=c=1$)- DQFT formulation emerges from a simple demand that a quantum state or a quantum field operator respects the discrete symmetries (Parity $\mathcal{P}$ and time reversal $\mathcal{T}$) of the background spacetime. Here we summarize the conceptual understanding of DQFT and direct the reader to Gaztañaga and Sravan Kumar (2024); Sravan Kumar and Marto (2023a, b) for more details. We start with the direct-sum Schrödinger equation
$i\frac{\partial|\Psi\rangle}{\partial t_{p}}=\begin{pmatrix}\hat{\mathbb{H}}&0\\ 0&-\hat{\mathbb{H}}\end{pmatrix}|\Psi\rangle$ (1)
@@
where $\mathbb{H}$ is the time-independent Hamiltonian (which is Hermitian) that is $\mathcal{P}\mathcal{T}$ symmetric. The quantum state $|\Psi\rangle$ is the direct-sum of two components
@@
where $|\Psi_{+}\rangle,\,|\Psi_{-}\rangle$ are the two components of the same state at parity conjugate points. The component $|\Psi_{+}\rangle$ is a positive energy state that evolves forward in time as $|\Psi_{+}\rangle_{t_{p}}=e^{-i\mathbb{H}t_{p}}|\Psi_{+}\rangle_{0}$ where the parametric time arrow $t_{p}:-\infty\to\infty$ while the $|\Psi_{-}\rangle$ can also be seen as positive energy state that evolves back in time as $|\Psi_{-}\rangle_{t_{p}}=e^{i\mathbb{H}t_{p}}|\Psi_{-}\rangle_{0}$ with the parametric time convention $t_{p}:+\infty\to-\infty$. One can notice here the purely anti-unitary character of time associated with the replacement $i\to-i$ Donoghue and Menezes (2019). The direct-sum Schrödinger equation does not have any arrow of time associated with it in comparison to the conventional Schrödinger equation. The direct-sum splitting of the quantum state (2) implies the Hilbert space splitting into a direct-sum of the two corresponding to parity conjugate points in position space: $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$. The position and momentum commutation relations now become double associated with parity conjugate position and momentum space operators denoted by subscripts +,+,\,-+ , - and their relations are
$\hat{p}_{+}=-i\frac{\partial}{\partial x_{+}}\quad,\quad\Big{[}\hat{x}_{+},\,\hat{p}_{+}\Big{]}$ $=i,\quad\Big{[}\hat{x}_{-},\,\hat{p}_{-}\Big{]}=-i\,$ (3)
$\hat{p}_{-}=+i\frac{\partial}{\partial x_{-}}\quad,\quad[\hat{x}_{+},\,\hat{x}_{-}]$ $=[\hat{p}_{+},\,\hat{p}_{-}]=0\,,$
@@
This framework only makes the wavefunction of the quantum system explicitly $\mathcal{P}\mathcal{T}$ symmetric that resonates with the assumed $\mathcal{P}\mathcal{T}$ symmetry of the physical system i.e., $[\mathbb{H},\,\mathcal{P}\mathcal{T}]=0$ and the results of standard quantum mechanics remains the same as shown with the example of harmonic oscillator worked out in Gaztañaga and Sravan Kumar (2024). The DQFT in Minkowski spacetime ($ds^{2}=-dt_{m}^{2}+d\textbf{x}^{2}$ which is $\mathcal{P}\mathcal{T}$ symmetric) is just built on the direct-sum Schrödinger equation, following how standard QFT is built. See Appendix A for more details and Fig.1 for an illustration.
Figure 1: The left-hand side depicts the picture of a light cone where a quantum field in standard QFT evolves from the past to the future while the field operators commute for spacelike distances $\left[\hat{\phi}\left(x\right),\ \hat{\phi}\left(y\right)\right]=0,\,\left(x-y\right)^{2}>0$ which is often called the causality condition. The right-hand side light cone represents how a quantum field in DQFT has two direct-sum components with opposite time evolutions at parity conjugate regions satisfying an additional new causality condition $\left[\hat{\phi}_{+}\left(x\right),\,\hat{\phi}_{-}\left(-x\right)\right]=0$.
@@
DQFT in curved spacetime, gravitational horizons unitarity and information-loss paradox ($G=\hbar=c=1$)- The de Sitter (dS) spacetime is a perfect testing ground for exploring the theory of QFTCS because of its maximally symmetric aspect and it is also relevant for unlocking the nature of inflationary quantum fluctuations. Furthermore, dS spacetime has the closest resemblance with black holes where the most important conundrums in quantum physics emerge such as the loss of unitarity and information-loss paradox.
We start with the dS metric in the flat Friedman-Lemaître-Robertson-Walker (FLRW) coordinates
@@
where the scale factor $a=e^{Ht}$ is the clock that determines the expansion of the Universe, $H=\frac{\dot{a}}{a}$ is the Hubble parameter and $\tau=\int\frac{dt}{a}=-\frac{1}{aH}$ is the conformal time. If one has to place quantum theory in dS spacetime (4) the first and foremost things to be understood are the parity and time reversal operations. The usual conception of understanding the expansion of the Universe is to restrict to $\tau<0$. But the metric carries a symmetry
@@
and restricting to $\tau<0$ before quantization of a classical field (as it is usually followed in standard QFTCS) means throwing away the symmetry by hand (5). The expanding Universe can be realized in two ways:
@@
This would explain why in the Wheeler-de Witt equation the cosmic time does not explicitly appear, rather the scale factor is the clock that runs with expanding Universe Kiefer (2007). Note that $H\to-H$ does not change at all any properties of dS space. For example, curvature invariants such as Ricci scalar $R=12H^{2}$ remain the same under discrete sign flip of H. The equation (6) illustrates two possible time realizations for an expanding universe. By convention, choosing $\tau>0$ or $\tau<0$ results in information loss beyond the horizon. Schrödinger in his famous monograph of 1956 Erwin Schrodinger (1956) rejected the idea of two Universes because it would lead to a conundrum of ignorance of information beyond the horizon which in modern language pure states evolving into mixed states which leads the violation of unitarity and information-loss Gibbons and Hawking (1977); Parikh et al. (2003).
@@
Similar to the Minkowski space (see Appendix A), the quantum fields in the dS space are created in a direct-sum vacuum $|0_{T}\rangle_{\rm dS}=|0_{+}\rangle_{\rm dS}\oplus|0_{-}\rangle_{\rm dS}$ corresponding to the direct-sum Fock space $\mathcal{F}_{\rm dS}=\mathcal{F}_{\rm dS+}\oplus\mathcal{F}_{\rm dS-}$. This implies, that everywhere in dS space a quantum field is tied by time forward and backward evolutions at the parity conjugate points, see Appendix B. At any moment of dS expansion, the parity-conjugate points on the comoving horizon $r_{H}=|\frac{1}{aH}|$ are space-like separated for an imaginary classical observer as shown in Fig. 2.
@@
Figure 2: This figure depicts the concept of unitarity that can be novelly achieved in de Sitter spacetime (in flat FLRW coordinates) with DQFT. Given the particle horizon ($r_{H}$) at any moment of de Sitter expansion the above picture depicts how DQFT formulation projects back all information of the states outside the horizon back inside at the antipodal point of the physical space. This implies all the information that any imaginary classical observer perceives is within the horizon and pure states cannot evolve into mixed states.
Figure 3: This picture depicts the conformal (not) Penrose diagram of static dS spacetime (8) where we depict direct-sum quantum states $|\phi_{+}\rangle\oplus|\phi_{-}\rangle$ entering and exiting the horizon with opposite time evolutions. This is not a Penrose diagram because each point represents half of the $S_{2}$ sphere in contrast to the usual Penrose diagram. This diagram is to represent quantum physics in dS space rather than classical physics for which the usual Penrose diagram is used.
@@
Note that by DQFT construction, the direct-sum quantum states $|\phi_{\pm}\rangle=\hat{d}_{(\pm)\,\textbf{k}}^{\dagger}|0_{\pm}\rangle_{\rm dS}$ that follow from (20) are the opposite time evolving positive energy states at the parity conjugate points and this holds for any imaginary classical observer. Furthermore, DQFT preserves the causality and locality too. A consequence of this is that any state that disappears beyond the horizon happens to be a direct-sum state that reappears at the antipodal point. This is because all the imaginary classical observers including those on the horizons of each other should observe the same direct-sum quantum states $|\phi_{+}\rangle\oplus|\phi_{-}\rangle$.
@@
This means the horizon acts as a $\mathcal{P}\mathcal{T}$ mirror which maps every information outside the horizon to the states inside the horizon. Fig. 2 gives the pictorial representation of this. This feature of DQFT implies any maximally entangled state (a pure state)
$|\phi\rangle\otimes|\psi\rangle=\frac{1}{\sqrt{2}}\left(|\phi_{+}\rangle\otimes|\psi_{+}\rangle\right)\oplus\frac{1}{\sqrt{2}}\left(|\phi_{-}\rangle\otimes|\psi_{-}\rangle\right)$ (7)
@@
remains within the horizon because no information is lost beyond the horizon. Every information of each of the entanglement pairs is mapped to within the horizon, thus the pure state will remain a pure state. In other words, DQFT forbids any entanglement beyond the gravitational horizons. DQFT constructs a direct-sum Hilbert space with the interior and exterior being separated by a superselection rule Wick et al. (1952); nLab authors (2023). Thus, there would be an observer complementarity. We can also understand the same from the point of view of dS space in static coordinates which reads as

### v2

# CMB parity asymmetry from unitary quantum gravitational physics

@@ Abstract
Longstanding anomalies in the Cosmic Microwave Background (CMB), including the low quadrupole moment and hemispherical power asymmetry, have recently been linked to an underlying parity asymmetry. We show here how this parity asymmetry naturally arises within a quantum framework that explicitly incorporates PT symmetry—parity (P) and time-reversal (T). This additional symmetry restores unitarity in quantum field theory (QFT) in curved spacetime. When applied to inflationary quantum fluctuations, this extended symmetry predicts parity asymmetry as a natural consequence of cosmic expansion, which inherently breaks time-reversal symmetry. Observational data strongly favor this unitary approach, with a likelihood ratio exceeding 650 times that of predictions from the standard inflationary framework. These results provide compelling evidence for the quantum gravitational origins of CMB parity asymmetry.
I Introduction
@@
Understanding and testing the interplay between gravity and quantum mechanics (QM) is the most important key to unraveling the new physics beyond the standard model of particle physics and Einstein’s general relativity (GR). The merge of GR and QM is an essential physics that can be unlocked at the length scales of gravitational horizons. For this, we perhaps need an in-depth understanding of how one can resolve an age-old conundrum in the concept of time between the GR and QM. The success of Quantum field theory (QFT) in Minkowski spacetime (Coleman (2018)) already teaches us several lessons in this regard. Indeed, QFT is the culmination of special relativity (SR) and QM, where the concept of time already enters into the conflict. In SR, the space and time are on equal footing, but in QM, spatial coordinates are elevated to operators while time is just a parameter due to its anti-unitary nature under the discrete transformation ($\mathcal{T}:t\to-t$). The SR and QM are merged by imposing the commutativity of operators corresponding to space-like distances. The underlying reason is that SR forbids any communication happening in space-like distances, and we respect it and achieve it by formulating new mathematical rules via second quantization. This is exactly the lesson one must keep in mind to understand quantum fields in curved spacetime. That is, we take lessons from classical physics and comply with them according to the rules of quantum physics. The quantum field theory in curved spacetime (QFTCS) is the first step in seeing GR and QM acting together besides the renormalizable ultra-violet complete quantum gravity that may be required at the Planck scales. Furthermore, QFTCS is a necessity in the context of understanding inflationary quantum fluctuations (Mukhanov and Chibisov (1981); Sasaki (1986); Albrecht et al. (1994)).
@@
In this paper, we address crucial questions related to the meaning of time reversal and parity inversions in the context of inflationary background spacetime to place quantum fields on it and uncover their true nature in the cosmic microwave background (CMB). We explain how direct-sum quantum field theory (DQFT), Kumar and Marto (2025, 2024); Gaztañaga et al. (2024)) is a necessity to solve the unitarity problem of QFTCS. The question of unitarity in QFTCS precedes the question of a completely renormalizable quantum gravity theory of Planck scales. In contrast to the widespread schools of thought that argue unitarity in QFTCS needs to be given up within GR and QM, we propose a new fundamental understanding of this issue and a new observational test using the parity asymmetry in the CMB. We find that the theory of inflationary quantum fluctuations with DQFT (which we call shortly DSI, which stands for ”Direct-Sum inflation”) fits up to 650 times better than the Standard Inflation (SI). This paper ultimately presents the picture of how a new theory of QFTCS emerged from fundamental questions of quantum gravitational physics that can potentially explain the observed CMB large-scale anomalies. The new results in this paper are further complemented by our more extended companion paper (Gaztañaga and Kumar (2024)).
II DQFT, in a nutshell ($\hbar=c=1$)
@@
DQFT formulation emerges from a simple reconstruction of a quantum state or a quantum field operator through the discrete (a)symmetries (parity $\mathcal{P}$ and time reversal $\mathcal{T}$) of the background spacetime. Here, we summarize the conceptual understanding of DQFT and direct the reader to Gaztañaga and Kumar (2024); Kumar and Marto (2025, 2024); Gaztañaga et al. (2024) for more details. Construction of QFT in Minkowski spacetime requires the definition of positive energy state and the arrow of time Donoghue and Menezes (2019). Here, we build quantum theory with two arrows of time using the direct-sum Schrödinger equation
$i\frac{\partial|\Psi\rangle}{\partial t_{p}}=\begin{pmatrix}\hat{\mathbb{H}}_{+}&0\\ 0&-\hat{\mathbb{H}}_{-}\end{pmatrix}|\Psi\rangle$ (1)
@@
where $\mathbb{H}=\mathbb{H}_{+}\oplus\mathbb{H}_{-}$ is the time-independent Hamiltonian operator (which is Hermitian) that is $\mathcal{P}\mathcal{T}$ symmetric. The quantum state $|\Psi\rangle$ is the direct-sum of two components
@@
where $|\Psi_{+}\rangle,\,|\Psi_{-}\rangle$ are the two components of the same state at parity conjugate points. The component $|\Psi_{+}\rangle$ is a positive energy state that evolves forward in time as $|\Psi_{+}\rangle_{t_{p}}=e^{-i\mathbb{E}t_{p}}|\Psi_{+}\rangle_{0}$ where the parametric time arrow $t_{p}:-\infty\to\infty$ while $|\Psi_{-}\rangle$ can also be seen as a positive energy state that evolves back in time as $|\Psi_{-}\rangle_{t_{p}}=e^{i\mathbb{E}t_{p}}|\Psi_{-}\rangle_{0}$ with the parametric time convention $t_{p}:+\infty\to-\infty$. One can notice here the purely antiunitary character of time associated with the replacement $i\to-i$ (Donoghue and Menezes (2019)). The direct-sum Schrödinger equation does not have any arrow of time associated with it compared to the conventional Schrödinger equation. The direct-sum splitting of the quantum state as in Eq.2 implies the Hilbert space ($\mathcal{H}$) splitting into a direct-sum of the two ($\mathcal{H}_{\pm}$) corresponding to parity conjugate points in position space: $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$. Here, the Hilbert spaces $\mathcal{H}_{\pm}$ are called the geometric superselection sectors (Gaztañaga et al. (2024); Kumar and Marto (2025)). The position and momentum commutation relations now become double associated with parity conjugate position and momentum space operators denoted by subscripts +,+,\,-+ , -, and their relations are
$\hat{p}_{+}=-i\frac{\partial}{\partial x_{+}}$ ; $\Big{[}\hat{x}_{+},\,\hat{p}_{+}\Big{]}=-\Big{[}\hat{x}_{-},\,\hat{p}_{-}\Big{]}=i\,$ (3)
$\hat{p}_{-}=+i\frac{\partial}{\partial x_{-}}$ ; $[\hat{x}_{+},\,\hat{x}_{-}]=[\hat{p}_{+},\,\hat{p}_{-}]=[\hat{x}_{\pm},\,\hat{p}_{\mp}]=0\,.$
@@
This framework only makes the wavefunction of the quantum system explicitly $\mathcal{P}\mathcal{T}$ symmetric that resonates with the assumed $\mathcal{P}\mathcal{T}$ symmetry of the physical system, i.e. $[\mathbb{H},\,\mathcal{P}\mathcal{T}]=0$ and the results of standard quantum mechanics remain the same as shown with the example of harmonic oscillator worked out in Gaztañaga and Kumar (2024) (see also further discussions and Fig.1 in Kumar and Marto (2025)). The DQFT in Minkowski spacetime ($ds^{2}=-dt_{m}^{2}+d\textbf{x}^{2}$ which is $\mathcal{P}\mathcal{T}$ symmetric) is just built on the direct-sum Schrödinger equation, following how the standard QFT is built on the definition of positive energy state and the causality condition (i.e., commutativity of operators for spacelike distances). See Appendix A for more details.
III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
@@
The de Sitter (dS) spacetime is a perfect testing ground for exploring the theory of QFTCS because of its maximally symmetric aspect, and it is also relevant for unlocking the nature of inflationary quantum fluctuations. Furthermore, a dS spacetime has the closest resemblance with black holes, where the most important conundrums in quantum physics emerge such as the loss of unitarity and the information-loss paradox. We start with the dS metric in the flat Friedman-Lemaître-Robertson-Walker (FLRW) coordinates
@@
where the scale factor $a=e^{Ht}$ is the clock that determines the expansion of the Universe, $H=\frac{\dot{a}}{a}$ is the Hubble parameter and $\tau=\int\frac{dt}{a}=\frac{1}{aH}$ is the conformal time. If one has to place quantum theory in dS spacetime Eq.4 the first and foremost things to be understood are the parity and time reversal operations. The usual conception of understanding the expansion of the Universe is to restrict ourselves to $\tau<0$. But the metric carries a symmetry
@@
Restricting to $\tau<0$ before quantization of a classical field (as is usually followed in standard QFTCS) means throwing away the symmetry in Eq.5 by hand. The expanding Universe can be realized in two ways:
@@
This would explain why in the Wheeler-de Witt equation the cosmic time does not explicitly appear, rather the scale factor is the clock that runs with expanding Universe (Kiefer (2007)). Note that $H\to-H$ does not change any properties of the dS space at all. For example, curvature invariants such as the Ricci scalar $R=12H^{2}$ remain the same under a discrete sign flip of H. Eq.6 illustrates two possible time realizations for an expanding universe. By convention, choosing $\tau>0$ or $\tau<0$ results in loss of information beyond the horizon. In his famous monograph of 1956, Schrodinger Erwin Schrodinger (1956) rejected the idea of two universes because it would lead to a conundrum of ignorance of information beyond the horizon, which in modern language means pure states evolving into mixed states which leads the violation of unitarity and information-loss (Gibbons and Hawking (1977); Parikh et al. (2003)).
@@
Similar to the Minkowski space (see Appendix A), the quantum fields in the dS space are described through a direct-sum vacuum $|0_{T}\rangle_{\rm dS}=|0_{+}\rangle_{\rm dS}\oplus|0_{-}\rangle_{\rm dS}$ corresponding to the direct-sum Fock space $\mathcal{F}_{\rm dS}=\mathcal{F}_{\rm dS+}\oplus\mathcal{F}_{\rm dS-}$. This implies that everywhere in dS space, a quantum field is tied by time forward and backward evolutions at the parity conjugate points, see Appendix B. At any moment of dS expansion, the parity-conjugate points on the comoving horizon $r_{H}=|\frac{1}{aH}|$ are space-like separated for an imaginary classical observer as shown in Fig. 1.
@@
Figure 1: This figure depicts the concept of unitarity that can be novelly achieved in de Sitter spacetime (in flat FLRW coordinates) with DQFT. Given the particle horizon ($r_{H}$) at any moment of de Sitter expansion, the above picture depicts how the DQFT formulation projects all information of the states outside the horizon back inside at the antipodal point of the physical space. This implies that all the information that any imaginary classical observer perceives is within the horizon and that pure states evolve into pure states. Blue arrows indicate accessible from inside the horizon, while dashed lines indicate that in the quasi-deSitter case the time evolution of $|\phi_{+}\rangle$ and $|\phi_{-}\rangle$ could be different, which is the origin of creating CMB parity asymmetry in the large scales.
@@
Note that by DQFT construction, the direct-sum quantum states $|\phi_{\pm}\rangle=\hat{d}_{(\pm)\,\textbf{k}}^{\dagger}|0_{\pm}\rangle_{\rm dS}$ that follow from Eq.26 are the opposite time evolving positive energy states at the parity conjugate points and this holds for any imaginary classical observer. Furthermore, in DQFT the causality and locality are preserved too. A consequence of this is that any state that disappears beyond the horizon happens to be the complimentary direct-sum (component) state that reappears at the antipodal point. This is because all the imaginary classical observers, including those on the horizons of each other, should observe the same direct-sum quantum state $|\phi_{+}\rangle\oplus|\phi_{-}\rangle$.
@@
This means the horizon acts as a $\mathcal{P}\mathcal{T}$ mirror which maps every information outside the horizon to the states inside the horizon. To see this, consider a pure state:
$|\psi\rangle=\sum_{m}d_{m}|\phi_{m}\rangle\,,$ (7)
where $|\phi_{m}\rangle$ are the eigenvectors, i.e. for any observable (an operator) $\hat{O}$ we have: $\hat{O}|\phi_{m}\rangle=\lambda_{m}|\phi_{m}\rangle$ where $\lambda_{m}$ are the corresponding eigenvalues. Each $\lambda_{m}$ will be observed in $|\psi\rangle$ with a probability: $P(\lambda_{m})=|d_{m}|^{2}$. Consider now the (deSitter) Hubble horizon $r_{H}=|\frac{1}{aH}|$ around an observer (as illustrated in Fig.1). In the standard approach, information is lost (a pure state becomes a mix state) because communication can not happen for space-like distances (i.e. outside the horizon) as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977), in analogy to what happens in Hawking radiation and the black hole information loss paradox 1974Natur.248...30H; Kumar and Marto (2024). This also implies the lost of unitarity in QFT on spacetimes with horizons. In the DQFT approach we have:
$|\psi\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}\sum d_{m}^{+}|\phi_{m}^{+}\rangle\\ \sum d_{m}^{-}|\phi_{m}^{-}\rangle\end{pmatrix}\equiv\frac{1}{\sqrt{2}}\begin{pmatrix}|\psi^{+}\rangle\\ |\psi^{-}\rangle\end{pmatrix}$ (8)
where $|\phi_{m}^{\pm}\rangle$ are the eigenvectors corresponding to each $\mathcal{P}\mathcal{T}$ separated Hilbert space. So that the observable probabilities and operators become:
$P(\lambda_{m}^{\pm})=|d_{m}^{\pm}|^{2}\quad\text{and}\quad\hat{O}=\begin{pmatrix}\hat{O}_{+}&0\\ 0&\hat{O}_{-}\end{pmatrix}$ (9)
Information is not lost in this case because there is not entanglement across the horizon. To see this more clearly, consider the case where $|\psi\rangle$ is a pure entangled system $|\psi_{AB}\rangle=|\psi_{A}\rangle\otimes|\psi_{B}\rangle$:
$|\psi_{AB}\rangle=\sum_{m,n}d_{mn}|\phi_{Am}\rangle\otimes|\phi_{Bn}\rangle\,,$ (10)
where $\otimes$ represents the tensor product of two Hilbert spaces $A$ and $B$ which are entangled (inseparable or correlated): $d_{mn}\neq d_{m}d_{n}$ with: $|\phi_{A}\rangle=\sum_{m}d_{m}|\phi_{Am}\rangle$ and $|\phi_{B}\rangle=\sum_{n}d_{n}|\phi_{Bn}\rangle$. For example, a pair of particles (across the horizon). In the DQFT approach, we have:
$|\psi_{AB}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}\sum d_{mn}^{+}|\phi_{Am}^{+}\rangle\otimes|\phi_{Bn}^{+}\rangle\\ \sum d_{mn}^{-}|\phi_{Am}^{-}\rangle\otimes|\phi_{Bn}^{-}\rangle\end{pmatrix}\equiv\frac{1}{\sqrt{2}}\begin{pmatrix}|\psi_{AB}^{+}\rangle\\ |\psi_{AB}^{-}\rangle\end{pmatrix}$ (11)
@@
where $|\phi_{Am}^{\pm}\rangle$ and $|\phi_{Bn}^{\pm}\rangle$ are the eigenvectors corresponding to each $\mathcal{P}\mathcal{T}$ separated Hilbert spaces, which remain within the horizon because no information is exchanged across $\mathcal{P}\mathcal{T}$ separated spaces. All information about the entanglement pairs is mapped within the horizon; thus, the pure state will remain pure. In other words, DQFT forbids any entanglement beyond the gravitational horizons. DQFT constructs a direct-sum Hilbert space with the interior and exterior separated by a geometric superselection rule (which is similar to the concept of superselection sectors in QFT Wick et al. (1952); nLab authors (2023)). Thus, there would be an observer complementarity (see Fig.3 in Kumar and Marto (2025)). We can also understand the same from the point of view of dS space in static coordinates, which reads as